AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Apr 5 01:28:15 2025
Author
*
:
Subject
*
:
<p>Run started 13:40</p> <p>Target density 4.3e11, Beam injection is ~1e7, Pressure in the ring is ~1e-10 mbar.</p> <p>Detectors have been moved IN with servo motors only (left arm pot = 36.5 mm, right arm pot =37.7 mm</p> <p>Time between injections is 12 seconds. Beam is in the ring for 3 seconds of the cycle.</p> <p>System wide checks OK</p> <p>Temperatures OK</p> <p>Leakage current plot attached -> temperature fluctuations (attachment 1)</p> <p>ADC data items stats (attachment 2)</p> <p>Merger and tape server OK</p> <p>All histograms and statistics zero'd</p> <p>14:17</p> <p>Cave access required. No beam</p> <p>15:24</p> <p>No beam. System wide checks ok. Merger ok.</p> <p>Attatchment 3: stats</p> <p>Attatchment 4: Temperatures generally okay, 3 were red before reload, one red after reload.</p> <p>Attatchment 5-8: Waveforms. Layout 7 and 8 look different from yesterday, consistnely different after several reloads.</p> <p>Attatchment 9: Leakage current</p> <p>Attatchment 10: LOWENERGYEXEY PLOTS</p> <p>17:12</p> <p>Tunning beam. System wide checks ok. Merger ok.</p> <p>Attatchment 11: stats</p> <p>Attatchment 12: Temperatures generally okay, 3 were red reloaed and stayed red but not so high.</p> <p>Attatchment 13-16: Waveforms. Were the same as previous check.</p> <p>Attatchment 17: Leakage current. They are tunning the beam so its a little bit changed.</p> <p>19:28</p> <p>Still hvae unstable noncooled beam, currently being worked on.</p> <p>Tunning beam. System wide checks ok. Merger ok.</p> <p>Attatchment 18: Temperatures generally okay, 3 were red reloaed and stayed red but not so high and were slighlty dicrease.</p> <p>Attatchment 19-22: Waveforms. Were the same as previous check.</p> <p>Attatchment 23: Leakage current. Slowly increasing from last check in, highest its been but conditions are not great at the moment.</p> <p>21:10</p> <p>We did not have any update about the beam. They still work on it.</p> <p>Tunning beam. System wide checks ok. Merger ok. Stats were ok.</p> <p>Attatchment 24: Temperatures generally okay, Same 3 were red and they increased a little bit. Nothing too concerning.</p> <p>Attatchment 25-28: Waveforms. Were the same as previous check.</p> <p>Attatchment 29: Leakage current. Contining to increase by small amount from last check.</p> <p> </p> <p> </p> <p> </p> <p> </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown